OpenStudy (anonymous):

The asteroid Hygiea orbits the Sun like the other planets. Its period is about 2030 days. What is its mean distance from the Sun?

OpenStudy (unklerhaukus):

Kepler's Law of Periods (K3) \[T^2=a^3\] where \(T\) is the period in earth years and \(a\) is a the mean distance in astronomical units

OpenStudy (unklerhaukus):

Solving for the distance \[a=T^{2/3}\] So convert 2030 days into years, T = 2030/365 and you result for a will be in astronomical units , if you want kilometres, 1 [AU] = 150 * 10^6 [km]